# Misspecification uncertainties in near-deterministic regression

### 02 Feb 2024 v1

@@ III Analysis of the generalization error in the near-deterministic limit
${\mathcal{G}}_{0}[\pi]=-\left\langle\ln\left|\int_{\mathcal{P}}\rho_{\mathcal{M}}({\mathcal{E}}({\mathbf{X}})|{\mathbf{X}},{\bm{\Theta}})\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}\right|\right\rangle_{\mathcal{X}}.$
Whilst ${\mathbf{Y}}={\mathcal{E}}({\mathbf{X}})$ is a unique maximum, as $\epsilon\to 0$ the marginal distribution (5) concentrates ${\bm{\Theta}}$ over pointwise optimal sets (POPS) of parameters
@@
Any model with a constant term will ensure ${\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})\neq\emptyset$, but under misspecification it clear that $\cap_{{\mathbf{X}}\in{\mathcal{X}}}{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})=\emptyset$. The conditional distribution $\rho_{\mathcal{M}}({\mathcal{E}}({\mathbf{X}})|{\mathbf{X}},{\bm{\Theta}})$ will sharpen as $\epsilon\to 0$, introducing a Jacobian into the integral over ${\bm{\Theta}}$. This can be contained in a pointwise mass function
${\rm m}_{\mathcal{E}}({\bm{\Theta}},{\mathbf{X}})$ $\equiv\int_{\mathcal{P}}\delta({\bm{\Theta}}-{\bm{\Theta}}^{\prime})\rho_{\mathcal{M}}({\mathcal{E}}({\mathbf{X}})|{\mathbf{X}},{\bm{\Theta}}^{\prime}){\rm d}{\bm{\Theta}}^{\prime}$
$=\int_{{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})}\frac{\delta({\bm{\Theta}}-{\bm{\Theta}}^{\prime})}{\sqrt{\|{\bm{\nabla}}_{\bm{\Theta}}{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}}^{\prime})\|}}{\rm d}{\bm{\Theta}}^{\prime}.$ (14)
This gives our first main result, a generalization error for deterministic surrogate models
@@
To avoid $1/\epsilon^{2}$ divergence in ${\mathcal{G}}_{0}[\pi]$ a valid $\pi({\bm{\Theta}})$ must satisfy the ‘covering’ constraint as $\epsilon\to 0$ of
@@ III.1 POPS-constrained loss minimizing ensemble ansatz
The covering constraint (16) shows, in the limit $\epsilon\rightarrow 0$, the minimizer of ${\mathcal{G}}_{0}$ concentrates onto the union of POPS for each training point. This suggests that a representation in terms of a discrete set of models is likely to capture its key features. Selecting some ${\bm{\Theta}}_{\mathbf{X}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})$ for each ${\mathbf{X}}\in{\mathcal{X}}$ and possibly some weighting ${\rm w}({\mathbf{X}})>0$, equation (16) is satisfied by the ensemble ansatz
@@
$\pi_{E}({\bm{\Theta}})\equiv\langle{\rm w}({\mathbf{X}})\delta({\bm{\Theta}}-{\bm{\Theta}}_{\mathbf{X}})\rangle_{\mathcal{X}}.$ (17)
It is simple to show the generalization error is finite and bounded by
@@
${\mathcal{G}}_{0}[\pi_{E}]\leq\left\langle\ln\left|{\rm w}({\mathbf{X}}){\rm m}_{\mathcal{E}}({\bm{\Theta}}_{\mathbf{X}},{\mathbf{X}})\right|\right\rangle_{\mathcal{X}},$ (18)
and the expected loss is ${\mathcal{L}}[\pi_{E}]=\langle{\rm w}({\mathbf{X}})L({\bm{\Theta}}_{\mathbf{X}})\rangle_{\mathcal{X}}$. However, whilst ${\mathcal{G}}_{0}[\pi_{E}]$ is finite, the vast majority of the pointwise fits ${\bm{\Theta}}_{\mathbf{X}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})$ will give very poor predictions away from ${\mathbf{X}}$. We therefore select each pointwise fit to minimize the (POPS-constrained) expected loss ${\mathcal{L}}[\pi_{E}]$, i.e.
${\bm{\Theta}}^{*}_{\mathbf{X}}\equiv\arg\min_{{\bm{\Theta}}_{\mathbf{X}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})}L({\bm{\Theta}}_{\mathbf{X}}).$ (19)
@@
The goal of this POPS-constrained loss-minimization is to ensure the pointwise fits will be as close as possible; for convex loss landscapes, this is measured by the Fisher information metric ${\bm{\mathcal{I}}}^{*}_{\mathcal{L}}$. By minimizing the ensemble spread, we aim to distribute an (albeit small) mass to other POPS, thereby reducing the generalization error. In addition, for specified regression problems $\pi_{E}({\bm{\Theta}})$ then reduces to the true minimizer $\pi^{*}_{\mathcal{L}}({\bm{\Theta}})=\delta({\bm{\Theta}}-{\bm{\Theta}}^{*}_{\mathcal{L}})$.
@@
When the loss landscape is not convex additional constraints will have to be imposed to minimize (and measure) the ensemble spread. More generally, POPS-constrained loss minimization with our ensemble ansatz is potentially sub-optimal in that some other ${\bm{\Theta}}_{\mathbf{X}}$, which are not loss minimizers, could be members of multiple POPS, i.e. ${\bm{\Theta}}_{\mathbf{X}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}}^{\prime})$, which is a priori beneficial to minimizing the generalization error. The loss minimization scheme (19) will minimize the average prediction error of each selected ${\bm{\Theta}}_{\mathbf{X}}$ across all other ${\mathbf{X}}^{\prime}\in{\mathcal{X}}$. Exploring this tradeoff (multiple POPS occupancy vs global prediction error) would entail a combinatorial scheme which our simple ensemble ansatz avoids. A key limitation of (19) is that many pointwise fits may be near-degenerate, raising the generalization error and, in general, underestimating the average deviation of any test point from the minimum loss prediction. However, the predicted max/min range of any test point is expected to be robust in the limit $N/P\to\infty$, as this is invariant to how we weight the ensemble.
@@ III.2 POPS-constrained ensemble for linear regression
We discuss how the epistemic covariance for small $N/P$ can be incorporated in a variational PAC-Bayes scheme in Appendix B. However, in the following, we concentrate on misspecification uncertainties in the limit $N/P\to\infty$, where the influence of $\bm{\Sigma}_{0}$ is negligble.
For linear models the POPS-constraint is simply ${\bm{\Theta}}_{\mathbf{X}}\cdot{\mathbf{F}}({\mathbf{X}})\equiv{\mathcal{E}}({\mathbf{X}})$. The loss minmizing pointwise fits ${\bm{\Theta}}^{*}_{{\mathbf{X}}}$ can be evaluated analytically via the efficient rank-one updates
@@
where ${\rm h}({\mathbf{X}})=[{\mathbf{F}}({\mathbf{X}})]^{\top}{\bf A}{\mathbf{F}}({\mathbf{X}})$ is the leverage, an outlier measure closely related to the distance of (Mahalanobis, 1936). The minimum loss solution is the leverage-weighted centroid of pointwise optima, ${\bm{\Theta}}^{*}_{\mathcal{L}}=\langle{\rm h}({\mathbf{X}}){\bm{\Theta}}^{*}_{{\mathbf{X}}}\rangle_{\mathcal{X}}$. Equation (21) is our third main result, a scheme to incorporate misspecification uncertainty into least-squares linear regression for minimal overhead, requiring only $\mathcal{O}(N)$ inner products over feature vectors. Implementation in least squares linear regression is trivial, as shown in the following python code, where A.T is the transpose of A and we assume a uniform weighting ${\rm w}({\bf X})=1/N$:
1import numpy as np
2def pointwise_optima(X,y):
3    # X.shape = (N,P), y.shape = (N,)
4    A = np.linalg.inv(np.dot(X.T,X))
5    # Minimum loss solution
6    Theta_L = np.dot(A,np.dot(X.T,y))
7    # Minimum loss errors
8    e = y - np.dot(X, Theta_L)
9    # Leverage of each point
10    h = ( X * np.dot(X, A) ).sum(1)
11    # Pointwise fits, shape = (P,N)
12    T = np.multiply(np.dot(X,A).T, e/h )
13    return T
An ensemble of test predictions can then be made simply by contracting with a constant tensor, a process which can easily be vectorized/parallelized. We now investigate how our POPS-constrained ensemble performs in a variety of high-dimensional regression tasks.
@@ IV.1 Comparison with direct minimization of $\mathcal{G}$
For low dimensional problems, it is possible to approximately minimize an empirical generalization error $\mathcal{G}$ over discrete model ensembles, using a finite aleatoric uncertainty ${\bm{\Sigma}}_{\mathcal{Y}}$ for regularization. With increasing ${\bm{\Sigma}}_{\mathcal{Y}}$ it is clear that $\mathcal{G}\to\mathcal{L}$ and thus we expect the minimum loss solution $\pi_{\mathcal{L}}({\bm{\Theta}})$ to emerge. However, at small ${\bm{\Sigma}}_{\mathcal{Y}}$ direct minimization typically fails, requiring careful parameter tuning and remaining at low $P<6$ to avoid vanishing gradients or other numerical issues. The POPS-ensemble, in contrast, can easily sample $P=\mathcal{O}(1000)$ or higher dimensional data.
Figure 1 shows a simple example fitting a one dimensional quadratic polynomial to a sinusoidal function, with the constant, linear and quadratic features giving $P=3$ model parameters. However, the features are highly correlated, simplifying the regression task. At large ${\bm{\Sigma}}_{\mathcal{Y}}$ the ensemble solution agrees closely with the minimum loss solution, whilst at small ${\bm{\Sigma}}_{\mathcal{Y}}$ the range of fits broadens, with the max/min range closely matching that of the POPS-ensemble. We also plot the result of Bayesian ridge regression as implemented in scikit-learn, showing that even at the ±$\pm 3\sigma$ range, predicted parameter uncertainties fail to capture the true test errors, which, as discussed above, will worsen with increasing $N/P$.
@@ IV.2 Test errors in medium-dimensional linear regression
We now look at higher/medium dimensional problems, $P=10-100$, where direct minimization of ${\mathcal{G}}$ is not practical. The simulation engines (see supplementary material) were constrained to be misspecified from the linear surrogate models by combining features in quadratic, cubic or sinusoidal functions or randomly sampling the linear feature coefficients from some predetermined subset of models. A noise term that simulates the presence of variability orthogonal to the chosen feature set was also used. In all cases, independent test and training data were generated, typically with a 10:90 test:train split. However, all conclusions remain robust to varying the precise nature of data generation and the classes of engine considered. Additional examples are provided in the supplementary material and appendix C.
@@ IV.2.1 POPS-ensemble prediction of test errors
Whilst the envelope bounds are highly informative, it is also desirable to predict test error distributions. As discussed above, the limitation of our POPS-ensemble is that the weighting of ensemble members is unlikely to be optimal, whilst the max/min range can be expected to be robust. As a result, a conservative estimate of the test error can be generated by uniformly resampling the max/min bounds for each test point to generate predicted error histograms. As shown in the top (right) of figure 3, this prediction from the POPS-ensemble results in excellent agreement with the test error ensemble, as confirmed across multiple engines in appendix A. Whilst a single (re)sample per test point is sufficient, it is possible to effectively bootstrap error histograms through repeated resampling.

### 09 Apr 2024 v2

@@ III Analysis of the generalization error in the near-deterministic limit
${\mathcal{G}}_{0}[\pi]$ $=-\left\langle\ln\left|\rho_{\mathcal{M}}({\mathcal{E}}({\mathbf{X}})|{\mathbf{X}})\right|\right\rangle_{\mathcal{X}}$ (12)
$=-\left\langle\ln\left|\int_{\mathcal{P}}\rho_{\mathcal{M}}({\mathcal{E}}({\mathbf{X}})|{\mathbf{X}},{\bm{\Theta}})\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}\right|\right\rangle_{\mathcal{X}}.$
Whilst ${\mathbf{Y}}={\mathcal{E}}({\mathbf{X}})$ is a unique maximum, to ensure $\min_{{\mathbf{X}}\in{\mathcal{X}}}\rho_{\mathcal{M}}({\mathcal{E}}({\mathbf{X}})|{\mathbf{X}})>0$ and thus avoid divergence in ${\mathcal{G}}_{0}[\pi]$ as $\epsilon\to 0$ we require $\pi({\bm{\Theta}})$ has mass in every pointwise optimal parameter set (POPS)
@@
Any model with a constant term will ensure ${\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})\neq\emptyset$, but under misspecification it clear that $\cap_{{\mathbf{X}}\in{\mathcal{X}}}{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})=\emptyset$.
To express the final generalization error, we introduce a pointwise mass function
${\rm m}_{\mathcal{E}}({\bm{\Theta}},{\mathbf{X}})$ $\equiv\int_{\mathcal{P}}\delta({\bm{\Theta}}-{\bm{\Theta}}^{\prime})\rho_{\mathcal{M}}({\mathcal{E}}({\mathbf{X}})|{\mathbf{X}},{\bm{\Theta}}^{\prime}){\rm d}{\bm{\Theta}}^{\prime},$
which gives our first main result, a generalization error for deterministic surrogate models of
@@
To avoid $1/\epsilon^{2}$ divergence in ${\mathcal{G}}_{0}[\pi]$ a valid $\pi({\bm{\Theta}})$ must satisfy the ‘POPS covering’ constraint as $\epsilon\to 0$ of
@@ III.1 POPS-constrained loss minimizing ensemble ansatz
The covering constraint (15) shows the minimizer of ${\mathcal{G}}_{0}$ must have mass in the POPS of each training point. This suggests that a representation in terms of a discrete set of models is likely to capture its key features. Selecting some set of pointwise fits $\mathcal{T}_{E}=\{{\bm{\Theta}}_{\mathbf{X}}\}_{{\mathbf{X}}\in{\mathcal{X}}}$, ${\bm{\Theta}}_{\mathbf{X}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})$ for each ${\mathbf{X}}\in{\mathcal{X}}$ and some weighting ${\rm w}_{E}({\mathbf{X}})>0$, $\langle{\rm w}_{E}({\mathbf{X}})\rangle_{\mathcal{X}}=1$, equation (15) is satisfied by the Gaussian ensemble ansatz
@@
$\pi_{E}({\bm{\Theta}})\equiv\langle{\rm w}_{E}({\mathbf{X}})\mathcal{N}({\bm{\Theta}}|{\bm{\Theta}}_{{\mathbf{X}}},\sigma^{2}\mathbb{I})\rangle_{\mathcal{X}},$ (16)
where $\sigma$ is fixed and assumed small. Defining mass function for the ensemble which measures the number of pointwise fits ${\bm{\Theta}}_{{\mathbf{X}}^{\prime}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})$
@@
$m_{E}({\mathbf{X}})\equiv\left\langle{\rm w}_{E}({\mathbf{X}}^{\prime}){\rm m}_{\mathcal{E}}({\bm{\Theta}}_{{\mathbf{X}}^{\prime}},{\mathbf{X}})\right\rangle_{{\mathcal{X}}},$ (17)
the generalization error of (16) reads, in the limit $\sigma\to 0$,
${\mathcal{G}}_{0}[\pi_{E}]=-\left\langle\ln|m_{E}({\mathbf{X}})|\right\rangle_{{\mathcal{X}}}.$ (18)
The above result shows to minimize ${\mathcal{G}}_{0}[\pi_{E}]$ we should select fits ${\bm{\Theta}}_{\mathbf{X}}\in\mathcal{T}_{E}$ to have maximum occupancy across all other POPS. In the worst case $\mathcal{T}_{E}\cap{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})={\bm{\Theta}}_{{\mathbf{X}}}$, $\forall{\mathbf{X}}\in{\mathcal{X}}$, ${\mathcal{G}}_{0}[\pi_{E}]$ still has a finite upper bound as $\sigma\to 0$ of
${\mathcal{G}}_{0}[\pi_{E}]\leq-\left\langle\ln|{\rm w}_{E}({\mathbf{X}})|\right\rangle_{{\mathcal{X}}}+\mathcal{S}_{{\mathcal{X}}},$ (19)
where $\mathcal{S}_{{\mathcal{X}}}=-\left\langle\ln|\rho({\mathbf{X}})|\right\rangle_{{\mathcal{X}}}$ is the entropy of the input distribution. In appendix D we show how each ${\bm{\Theta}}_{\mathbf{X}}\in\mathcal{T}_{E}$ can be determined in a variational setting at finite $\sigma$, using (16) in the PAC-Bayes loss bound (7) for the regression loss ${\mathcal{L}}[\pi_{E}]=\langle{\rm w}_{E}({\mathbf{X}})L({\bm{\Theta}}_{\mathbf{X}})\rangle_{\mathcal{X}}$, which as $P/N\to 0$, $\sigma\to 0$ reduces to POPS-constrained loss minimization
${\bm{\Theta}}^{*}_{\mathbf{X}}$ $\equiv\arg\min_{{\bm{\Theta}}_{\mathbf{X}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})}L({\bm{\Theta}}_{\mathbf{X}}),$ (20)
$\pi^{*}_{E}({\bm{\Theta}})$ $\equiv\langle{\rm w}_{E}({\mathbf{X}})\delta({\bm{\Theta}}-{\bm{\Theta}}^{*}_{{\mathbf{X}}})\rangle_{\mathcal{X}}.$ (21)
@@
The goal of this POPS-constrained loss-minimization is to ensure the pointwise fits will be as close as possible; for convex loss landscapes, this is measured by the Fisher information metric ${\bm{\mathcal{I}}}^{*}_{\mathcal{L}}$. By minimizing the ensemble spread, we aim to distribute an (albeit small) mass to other POPS, thereby reducing ${\mathcal{G}}_{0}$. In addition, for specified regression problems $\pi^{*}_{E}({\bm{\Theta}})$ then reduces to the true minimizer $\pi^{*}_{\mathcal{L}}({\bm{\Theta}})=\delta({\bm{\Theta}}-{\bm{\Theta}}^{*}_{\mathcal{L}})$.
@@
When the loss landscape is not convex additional constraints will have to be imposed to minimize (and measure) the ensemble spread. More generally, POPS-constrained loss minimization with our ensemble ansatz is potentially sub-optimal in that $\pi^{*}_{E}({\bm{\Theta}})$ chooses ${\bm{\Theta}}^{*}_{\mathbf{X}}$ to minimize the average prediction error across all other ${\mathbf{X}}^{\prime}\in{\mathcal{X}}$ but is not guaranteed to be a minimizer of (18). Exploring this tradeoff (multiple POPS occupancy vs global prediction error) would entail a combinatorial scheme which (21) avoids. One strategy, to be investigated in future work, would be to cluster the loss-minimizing pointwise fits, try to find POPS-intersections for each cluster, then replace all members of that cluster with a single POPS-intersecting fit. A further limitation of (21) is that many pointwise fits may be near-degenerate, raising the generalization error and, in general, underestimating the average deviation of any test point from the minimum loss prediction. However, the predicted max/min range of any test point is expected to be robust in the limit $N/P\to\infty$, as this is invariant to how we weight the ensemble.
@@ III.2 POPS-constrained ensemble for linear regression
We discuss how the epistemic covariance for small $N/P$ can be incorporated in a variational PAC-Bayes scheme in Appendix D. However, in the following, we concentrate on misspecification uncertainties in the limit $N/P\to\infty$, where the influence of $\bm{\Sigma}_{0}$ is negligible. For linear models the POPS-constraint is simply ${\bm{\Theta}}_{\mathbf{X}}\cdot{\mathbf{F}}({\mathbf{X}})\equiv{\mathcal{E}}({\mathbf{X}})$. The loss minimizing pointwise fits ${\bm{\Theta}}^{*}_{{\mathbf{X}}}$ can be evaluated analytically via the efficient rank-one updates
@@
where ${\rm h}({\mathbf{X}})=[{\mathbf{F}}({\mathbf{X}})]^{\top}{\bf A}{\mathbf{F}}({\mathbf{X}})$ is the leverage, an outlier measure closely related to the distance of (Mahalanobis, 1936). The minimum loss solution is the leverage-weighted centroid of pointwise optima, ${\bm{\Theta}}^{*}_{\mathcal{L}}=\langle{\rm h}({\mathbf{X}}){\bm{\Theta}}^{*}_{{\mathbf{X}}}\rangle_{\mathcal{X}}$. Equation (23) is our third main result, a scheme to incorporate misspecification uncertainty into least-squares linear regression for minimal overhead, requiring only $\mathcal{O}(N)$ inner products over feature vectors. Implementation in least squares linear regression is trivial, as shown in appendix F. We now investigate how our POPS-constrained ensemble performs in a variety of high-dimensional regression tasks.
@@ IV.1 Comparison with direct minimization of $\mathcal{G}$
For low dimensional problems, it is possible to approximately minimize an empirical generalization error $\mathcal{G}$ over discrete model ensembles, using a finite aleatoric uncertainty ${\bm{\Sigma}}_{\mathcal{Y}}$ for regularization. With increasing ${\bm{\Sigma}}_{\mathcal{Y}}$ it is clear that $\mathcal{G}\to\mathcal{L}$ and thus we expect the minimum loss solution $\pi_{\mathcal{L}}({\bm{\Theta}})$ to emerge. We also compare these appraoches to Bayesian ridge regression, corresponding to minimizing a first-order PAC-Bayes loss bound, and an ensemble approach from citepmasegosa2020learning, section C.3, which corresponds to minimizing a second-order PAC-Bayes loss bound, whose loss function, as for $\mathcal{G}_{E}$, also requires exponentiating model errors. At small ${\bm{\Sigma}}_{\mathcal{Y}}$ direct minimization of $\mathcal{G}_{E}$ and the Masegosa ensemble bound typically fails, requiring careful parameter tuning and remaining at low $P<6$ to avoid vanishing gradients or other numerical issues. The POPS-ensemble, in contrast, can easily sample $P=\mathcal{O}(1000)$ or higher dimensional data.
Figure 1 shows a simple example fitting a one dimensional quadratic polynomial to a sinusoidal function, with the constant, linear and quadratic features giving $P=3$ model parameters. However, the features are highly correlated, simplifying the regression task. We show results for ${\bm{\Sigma}}_{\mathcal{Y}}=1/3$, at the lower limit for which minimization converges for all methods. As expected from section II.3 Bayesian ridge regression significantly underestimates test errors even at the ±$\pm 3\sigma$ level, which as discussed above will worsen with increasing $N/P$. We find similar behaviour for the Masegosa ensemble, which we note is primarily designed for cases with finite aleatoric uncertainty. Direct minimization of $\mathcal{G}_{E}$ gives better bounding of test errors, though still with notable underestimates in high-density regions, though bounnds became tighter with lower ${\bm{\Sigma}}_{\mathcal{Y}}$ (whilst stable) However, the POPS-constrained ensemble gives almost perfect bounding of test errors, being stable for arbitrarly small ${\bm{\Sigma}}_{\mathcal{Y}}$ values.
@@ IV.2 Test errors in medium-dimensional linear regression
We now look at higher/medium dimensional problems, $P=10-100$, where direct minimization of ${\mathcal{G}}$ is not practical. The simulation engines (see supplementary material) were constrained to be misspecified from the linear surrogate models by combining features in quadratic, cubic or sinusoidal functions or randomly sampling the linear feature coefficients from some predetermined subset of models. A noise term that simulates the presence of variability orthogonal to the chosen feature set was also used. In all cases, independent test and training data were generated, typically with a 10:90 test:train split. However, all conclusions remain robust to varying the precise nature of data generation and the classes of engine considered. Additional examples are provided in the supplementary material and appendix G.
@@ IV.2.1 POPS-ensemble prediction of test errors
Whilst the envelope bounds are highly informative, it is also desirable to predict test error distributions. As discussed above, the limitation of our POPS-ensemble is that the weighting of ensemble members is unlikely to be optimal, whilst the max/min range can be expected to be robust. As a result, a conservative estimate of the test error can be generated by uniformly resampling the max/min bounds for each test point to generate predicted error histograms. As shown in the top (right) of figure 3, this prediction from the POPS-ensemble results in excellent agreement with the test error ensemble, as confirmed across multiple engines in appendix B. Whilst a single (re)sample per test point is sufficient, it is possible to effectively bootstrap error histograms through repeated resampling.